# FinRAGBench-V: A Benchmark for Multimodal RAG with Visual Citation in the Financial Domain

### v1

@@ Abstract
Suifeng Zhao1, Zhuoran Jin2, Sujian Li3*, Jun Gao1thanks: Corresponding Authors 1Key Laboratory of High Confidence Software Technologies, School of Computer Sciences, Peking University 2School of Artificial Intelligence, University of Chinese Academy of Sciences 3State Key Laboratory of Multimedia Information Processing, School of Computer Sciences, Peking University zhaosuifeng826@gmail.com, zhuoran.jin@nlpr.ia.ac.cn, {lisujian,gaojun}@pku.edu.cn
@@ 1 Introduction
However, existing financial RAG efforts, such as FinQA Chen et al. (2021) and OmniEval Wang et al. (2024b), predominantly focus on text-only RAG, which may lose critical information when converting multimodal documents into plain text. As a result, they frequently fail to answer questions accurately, as shown in Figure 1 (a). Although MME-Finance (Gan et al., 2024) introduces a multimodal reasoning benchmark, it relies mostly on isolated screenshots and lacks retrieval support. Consequently, it falls short of reflecting the complexity of real-world financial scenarios, where answering questions often requires diverse data sources and heterogeneous data types. Furthermore, given the critical importance of precision in finance, RAG systems must ensure not only accuracy responses but also their traceability and verifiability, yet most existing benchmarks overlook these needs. Thus, designing a more comprehensive benchmark for multimodal RAG in finance is imperative.
In this work, we propose FinRAGBench-V, a multimodal RAG benchmark tailored for finance, featuring grounded visual citation. This benchmark effectively integrates multimodal data and provides visual citations to ensure traceability, as shown in Figure 1 (b). Specifically, we construct a large-scale retrieval corpus from diverse real-world financial sources, comprising 60,780 Chinese pages from 1,104 documents and 51,219 English pages from 1,105 documents, including research reports, financial statements, prospectuses, etc. In addition, we develop a high-quality financial question-answering (QA) dataset using GPT-4o assistance with manual verification. The dataset consists of 855 Chinese and 539 English QA pairs, covering a wide range of distinctive financial tasks, with questions categorized by data heterogeneity, including text, charts, and tables, and reasoning type, such as time-sensitive reasoning, numerical calculations, multi-page reasoning, etc.
@@
We conduct extensive experiments and evaluations on FinRAGBench-V. For retrieval, we conduct experiments using four textual retrievers, such as Jina-ColBERT-V2 Jha et al. (2024), and five Multilingual-E5-large Wang et al. (2024a); and multimodal ones, such as ColQwen2 Faysse et al. (2024), GME-Qwen2-VL-2B Zhang et al. (2024b), and DSE-QWen2-2b-MRL-V1 Ma et al. (2024a). For generation and citation, we employ seven proprietary Multimodal Large Language Models (MLLMs), such as GPT-4o, GPT-4V, and Gemini-2.0-Flash, and six open-source ones, such as Qwen2.5-VL-72B-Instruct and MiniCPM-o-2.6.
@@
OmniEval Wang et al. (2024b) Finance
@@ 2 Related Work
Retrieval-Augmented Generation (RAG) has gained significant attention as an effective method of leveraging retrieval mechanisms to provide external knowledge to LLMs’ generation (Gao et al., 2023b; Lewis et al., 2020; Huang et al., 2023; Chen et al., 2024b; Friel et al., 2024; Saad-Falcon et al., 2024). In the financial domain, where charts and graphs are essential, text-only RAG benchmarks often overlook critical information (Chen et al., 2021; Wang et al., 2024b), highlighting the need for a multimodal RAG benchmark. Recent efforts on financial multimodal benchmarks exhibit several limitations, as summarized in Table 1. EvoChart (Huang et al., 2025) focuses solely on chart-based questions, lacking integration with textual and tabular information. Cho et al. (2024) and Suri et al. (2024) utilize real-world PDFs but support only limited question types. MME-Finance (Gan et al., 2024) provides diverse financial questions, yet its reliance on isolated chart screenshots hinders document-level retrieval and fails to reflect the complexity of financial data.
@@ 4.2.2 Quality Inspection
During the selection and annotation process, we adhere to several key principles to ensure the high quality and consistency of the dataset: examining the clarity of the questions and their correct categorization, verifying the accuracy of answers, and checking whether the page sources for multi-page queries are properly identified. Based on these criteria, we carefully filter and refine the original 11,328 generated QA pairs, and ultimately obtaining a total of 1,394 pairs, consisting of 855 Chinese entries and 539 English entries. The statistics of each category are shown in Figure 3, the lengths statistics of the dataset are shown in Table 3.
@@ 5.2 Generation with Visual Citation
During the generation stage, based on the retrieval result, the generator model $M$ is tasked with producing textual answer $a$ accompanied by visual citations $C$, given the query $q$. To enable the simultaneous generation of both answers and citations, we follow the visual citation method used in VISA Ma et al. (2024b). Specifically, we input both the question $q$ and the top-k relevant pages $\{p_{1},p_{2},...,p_{k}\}$ into the generator $M$, instructing it to generate the answer $a$ while simultaneously producing both page-level and block-level citations. Each citation is denoted as $c_{i}=(p_{i},\{b_{i1},b_{i2},...,b_{ij},...\})$, where the page-level citation $p_{i}$ refers to the reference page, $\{b_{i1},b_{i2},...,b_{ij},...\}$ represents the block-level citations, indicating the specific regions of the answer within the page. Each block-level citation $b_{ij}$ is represented as a set of coordinates, i.e., $b_{ij}=\left[x_{1},y_{1},x_{2},y_{2}\right]$, where $(x_{1},y_{1})$ and $(x_{2},y_{2})$ denote the coordinates of the top-left corner and bottom-right corner of $b_{ij}$, respectively. Detailed output format is in Appendix A.
@@ 6.3 Citation Quality
Recall evaluates whether the cited images are sufficient to support the answer. If the union of the citation set $C=\{c_{1},c_{2},...,c_{n}\}$ of an answer $a$ sufficiently support $a$, the recall is assigned $1$; otherwise, it is assigned 00, defined in Equation 1:
@@
It should be noted that the precision of each citation is evaluated only when the recall of the citation set it belongs to is $1$; otherwise, i is set to 00.
@@ 7.1 Basic Settings
In the generation phase, we conduct experiments on both proprietary and open-source MLLMs. The former consists of o4-mini, GPT-4o, GPT-4V, GPT-4o-mini, Gemini-1.5-Flash, Gemini-2.0-Flash, and Claude-3.5-Sonnet-20240620; while the later comprises Qwen2-VL-72B-Instruct, Qwen2.5-VL-7B-Instruct, Qwen2.5-VL-32B-Instruct, Qwen2.5-VL-72B-Instruct, Phi-3.5-vision-instruct, and MiniCPM-o-2.6. The prompt for generation is in Appendix A, more details are in Appendix F.
@@ 7.3 Consistency between Automatic Citation Evaluation with Human Evaluations
To complement IoU, we use human ratings of the predicted citations on a 0–5 scale, considering factors such as page and block relevance, offset from ground truth, and the inclusion of redundant or irrelevant content. This provides a more nuanced and semantically meaningful assessment of citation quality. The guideline for rating is shown in Appendix G.
@@ Appendix B Examples of Six Real-World Data Sources of Retrieval Corpus
Appendix C Examples of Seven Categories of QA Dataset
In this section, we provide an example for each category of questions, shown in Table 13, 14 C.2, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26.
C.1 Text Inference:
@@ C.7 Multi-page Queries
Appendix D Example for Visual Citation and the Two Evaluation Methods
@@
Appendix E Case Study
E.1 Error Case Study Based on Different Stages in RGenCite
@@ E.2 Error Case Study Based on Typical Task Types in the Financial Domain
Appendix F Resource Usage
Appendix G Annotation guideline for the Rating-based Human Evaluation.
This section demonstrates the annotation guideline for the rating-based human evaluation in Table 27.
Appendix H Potential Risks

### v2

@@ Abstract
Suifeng Zhao1, Zhuoran Jin2, Sujian Li3*, Jun Gao1thanks: Corresponding authors 1Key Laboratory of High Confidence Software Technologies, CS, Peking University, China 2School of Artificial Intelligence, University of Chinese Academy of Sciences 3State Key Laboratory of Multimedia Information Processing, School of Computer Sciences, Peking University sfzhao25@stu.pku.edu.cn, zhuoran.jin@nlpr.ia.ac.cn, {lisujian,gaojun}@pku.edu.cn [Uncaptioned image] Dataset: https://huggingface.co/datasets/zhaosuifeng/FinRAGBench-V [Uncaptioned image] Code: https://github.com/zhaosuifeng/FinRAGBench-V
@@ 1 Introduction
However, existing financial RAG efforts, such as FinQA Chen et al. (2021) and OmniEval Wang et al. (2024c), predominantly focus on text-only RAG, which may lose critical information when converting multimodal documents into plain text. As a result, they frequently fail to answer questions accurately, as shown in Figure 1 (a). Although MME-Finance (Gan et al., 2024) introduces a multimodal reasoning benchmark, it relies mostly on isolated screenshots and lacks retrieval support. Consequently, it falls short of reflecting the complexity of real-world financial scenarios, where answering questions often requires diverse data sources and heterogeneous data types. Furthermore, given the critical importance of precision in finance, RAG systems must ensure not only accuracy responses but also their traceability and verifiability, yet most existing benchmarks overlook these needs. Thus, designing a more comprehensive benchmark for multimodal RAG in finance is imperative.
In this work, we propose FinRAGBench-V, a multimodal RAG benchmark tailored for finance, featuring grounded visual citation. This benchmark effectively integrates multimodal data and provides visual citations to ensure traceability, as shown in Figure 1 (b). Specifically, we construct a large-scale retrieval corpus from diverse real-world financial sources, comprising 60,780 Chinese pages from 1,104 documents and 51,219 English pages from 1,105 documents, including research reports, financial statements, prospectuses, etc. In addition, we develop a high-quality financial question-answering (QA) dataset using GPT-4o OpenAI (2023) assistance with manual verification. The dataset consists of 855 Chinese and 539 English QA pairs, covering a wide range of distinctive financial tasks, with questions categorized by data heterogeneity, including text, charts, and tables, and reasoning type, such as time-sensitive reasoning, numerical calculations, multi-page reasoning, etc.
@@
We conduct extensive experiments and evaluations on FinRAGBench-V. For retrieval, we conduct experiments using four textual retrievers, such as Jina-ColBERT-V2 Jha et al. (2024), and five Multilingual-E5-large Wang et al. (2024a); and multimodal ones, such as ColQwen2 Faysse et al. (2024), GME-Qwen2-VL-2B Zhang et al. (2024b), and DSE-QWen2-2b-MRL-V1 Ma et al. (2024a). For generation and citation, we employ seven proprietary Multimodal Large Language Models (MLLMs), such as GPT-4o OpenAI (2023), GPT-4V, and Gemini-2.0-Flash Comanici et al. (2025), and six open-source ones, such as Qwen2.5-VL-72B-Instruct Wang et al. (2024b) and MiniCPM-o-2.6 Yao et al. (2024).
@@
OmniEval Wang et al. (2024c) Finance
@@ 2 Related Work
Retrieval-Augmented Generation (RAG) has gained significant attention as an effective method of leveraging retrieval mechanisms to provide external knowledge to LLMs’ generation (Gao et al., 2023b; Lewis et al., 2020; Huang et al., 2023; Chen et al., 2024b; Friel et al., 2024; Saad-Falcon et al., 2024). In the financial domain, where charts and graphs are essential, text-only RAG benchmarks often overlook critical information (Chen et al., 2021; Wang et al., 2024c), highlighting the need for a multimodal RAG benchmark. Recent efforts on financial multimodal benchmarks exhibit several limitations, as summarized in Table 1. EvoChart (Huang et al., 2025) focuses solely on chart-based questions, lacking integration with textual and tabular information. Cho et al. (2024) and Suri et al. (2024) utilize real-world PDFs but support only limited question types. MME-Finance (Gan et al., 2024) provides diverse financial questions, yet its reliance on isolated chart screenshots hinders document-level retrieval and fails to reflect the complexity of financial data.
@@ 4.2.2 Quality Inspection
During the selection and annotation process, we adhere to several key principles to ensure the high quality and consistency of the dataset: examining the clarity of the questions and their correct categorization, verifying the accuracy of answers, and checking whether the page sources for multi-page queries are properly identified. The detailed annotation guideline is shown in Table 27 of Appendix F. Based on these criteria, we carefully filter and refine the original 11,328 generated QA pairs, and ultimately obtaining a total of 1,394 pairs, consisting of 855 Chinese entries and 539 English entries. The statistics of each category are shown in Figure 3, the lengths statistics of the dataset are shown in Table 3.
@@ 5.2 Generation with Visual Citation
[…]
@@ 6.3 Citation Quality
Recall evaluates whether the cited images are sufficient to support the answer. If the union of the citation set $C=\{c_{1},c_{2},...,c_{n}\}$ of an answer $a$ sufficiently support $a$, the recall is assigned $1$; otherwise, it is assigned 0, defined in Equation 1:
@@
It should be noted that the precision of each citation is evaluated only when the recall of the citation set it belongs to is $1$; otherwise, i is set to 0.
@@ 7.1 Basic Settings
In the generation phase, we conduct experiments on both proprietary and open-source MLLMs. The former consists of o4-mini, GPT-4o OpenAI (2023), GPT-4V, GPT-4o-mini, Gemini-1.5-Flash Reid et al. (2024), Gemini-2.0-Flash Comanici et al. (2025), and Claude-3.5-Sonnet-20240620 Anthropic (2024); while the later comprises Qwen2-VL-72B-Instruct Wang et al. (2024b), Qwen2.5-VL-7B-Instruct, Qwen2.5-VL-32B-Instruct, Qwen2.5-VL-72B-Instruct, Phi-3.5-vision-instruct Abdin et al. (2024), and MiniCPM-o-2.6 Yao et al. (2024). The prompt for generation is in Appendix A, more details are in Appendix G.
@@ 7.3 Consistency between Automatic Citation Evaluation with Human Evaluations
To complement IoU, we use human ratings of the predicted citations on a 0–5 scale, considering factors such as page and block relevance, offset from ground truth, and the inclusion of redundant or irrelevant content. This provides a more nuanced and semantically meaningful assessment of citation quality. The guideline for rating is shown in Table 28 of Appendix F.
@@ 8 Conclusion
Acknowledgement
We thank the anonymous reviewers for their helpful comments on this paper. This work is supported by the National Natural Science Foundation of China (No. 62476010) and National Natural Science Foundation of China (No. 62272008).
@@ Appendix B Examples of Six Real-World Data Sources of Retrieval Corpus
Appendix C Examples of Seven Categories of QA Dataset
In this section, we provide an example for each category of questions, shown in Table 13, 14 H, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26.
C.1 Text Inference:
C.2 Chart-Information Extraction
C.3 Chart-Numerical Calculations
C.4 Chart-Time Sensitive
C.5 Table-Numerical Calculations
C.6 Table-Comparison and Sorting
C.7 Multi-page Queries
Appendix D Example for Visual Citation and the Two Evaluation Methods
Appendix E Case Study
E.1 Error Case Study Based on Different Stages in RGenCite
Recognizing Candlestick Charts.
Dealing with Complex Financial Table.
Dealing with Multi-page Questions.
Appendix F Annotation guidelines.
This section demonstrates the annotation guidelines. The annotation guidelines for constructing the QA dataset in shown Table 27 and the guideline for rating-based human evaluation for visual citation is in Table 28.
Appendix G Resource Usage
Appendix H Potential Risks
@@
Annotation Guideline for QA Pairs Verification
GUIDELINE: Please verify the QA pairs produced by GPT-4o. For each sample, you may choose to retain, revise, or discard it. Your decision should be based on the following four criteria:
Verification Criteria:
1. Query Clarity: The query should be specific and unambiguous, targeting a particular topic in a document, and avoiding vague or overly general questions.
2. Answer Correctness: The answer must be factually correct and directly supported by the visual content. It should not include hallucinations or inferred information beyond what is presented. If calculation is involved, the answer should be accurate.
3. Category Appropriateness: The question should match its assigned category (e.g., table-numerical calculations). Mislabelled or ambiguous categories should lead to revision or rejection.
4. Correctness of Multi-Page Sources: For multi-page queries, if the answer is derived from multiple pages, all referenced page sources must be accurately identified.
Decision Rule: Retain the data if all criteria are met, revise it if there are minor issues (e.g., unclear query, incorrect category), and discard it if there are major errors or cannot be fixed reliably.
Table 27: Annotation guideline for QA Pairs Verification